# Network Slicing Resource Management in Uplink User-Centric Cell-Free Massive MIMO Systems

### v1

@@ V-3 Comparison of Computational Complexity Performance
Fig. 4 presents the average runtime comparison between the Hybrid and Proposed schemes. As discussed in Section V-1, the Proposed scheme exhibits a substantially lower computational complexity, reduced by up to two orders in $K$ compared to the Hybrid scheme. When the CVX solver operates successfully, the Proposed scheme achieves up to $97\%$ lower runtime for lower UE densities. However, at higher UE densities, both schemes exhibit nearly identical runtimes, implying that the Hybrid scheme’s CVX optimization is bypassed due to feasibility violations under Proposition 2. In contrast, the Proposed scheme maintains consistent low complexity across all loads. Despite its lightweight computation, the Proposed scheme maintains superior QoS satisfaction and only a modest loss in weighted sum-rate compared to the Hybrid scheme. This highlights its overall computational efficiency and practicality for QoS-critical, resource-constrained large-scale deployments.
@@ VI Conclusion
In this paper, we present a computationally efficient framework for jointly optimizing per-UE bandwidth allocation and UE–AP association assignment in the UL of a NS-enabled CF mMIMO system, aiming to maximize the weighted sum-rate while satisfying heterogeneous QoS requirements for eMBB and URLLC UEs. Owing to the NP-hard nature of the formulated problem, it is decomposed into two tractable sub-problems, each addressed by an efficient heuristic scheme designed to ensure fair QoS distribution even under minimum bandwidth violation scenarios. Numerical results demonstrate that the proposed scheme achieves a favorable balance between computational efficiency, weighted sum-rate performance, and QoS satisfaction, making it a robust choice for QoS-critical network deployments. While this work focuses on maintaining service for all UEs through fair QoS distribution under resource scarcity, future extensions could explore admission control strategies to further enhance strict QoS guarantees for latency-critical applications.
@@ A-A Proof of Theorem 1
Problem P0 contains $KM$ binary association variables $a_{k,m}\in{0,1}$ and $K$ continuous bandwidth allocation variables $b_{k}\in\mathbb{R}+$, so it is a mixed-integer program. Nonconvexity follows from the objective’s dependence on the quadratic/bilinear terms (e.g., $(\sum_{m}a_{k,m}\gamma_{k,m})^{2}$ and $R_{k}=\left(1-\frac{\tau_{p}}{\tau_{c}}\right)b_{k}\log_{2}(1+\mathrm{SINR}_{k}(\mathbf{A}))$), together with the discrete feasible set $\{0,1\}^{KM}$.
For hardness, consider a restricted instance where bandwidths $b_{k}$ are fixed and QoS constraints are removed. Further, choose channel parameters (or assume negligible inter-UE interference) so that each $\mathrm{SINR}_{k}$ depends only on the AP assigned to UE $k$; then the objective reduces to maximizing $\sum_{k}w_{k}\log_{2}(1+\mathrm{SINR}_{k}(\mathbf{A}))$ that depend only on the assignment variables. Under this construction P0 subsumes the generalized assignment problem (GAP): assigning items (UEs) to bins (APs) with capacity constraints and profits, which is NP-hard [9]. Hence, by reduction from GAP, P0 is NP-hard.

### v2

@@ V-3 Comparison of Computational Complexity Performance
Fig. 4 presents the average runtime comparison between the Hybrid and Proposed schemes. As discussed in Section V-1, the Proposed scheme exhibits a substantially lower computational complexity, reduced by up to two orders in $K$ compared to the Hybrid scheme. When the CVX solver operates successfully, the Proposed scheme achieves up to $97\%$ lower runtime for lower UE densities. However, at higher UE densities, both schemes exhibit nearly identical runtimes, implying that the Hybrid scheme’s CVX optimization is bypassed due to feasibility violations under Proposition 2. In contrast, the Proposed scheme maintains consistent low complexity across all loads. Despite its lightweight computation, the Proposed scheme maintains superior QoS satisfaction and only a modest loss in weighted sum-rate compared to the Hybrid scheme. This highlights the Proposed scheme’s overall computational efficiency and practicality for QoS-critical, resource-constrained large-scale deployments.
@@ VI Conclusion
In this paper, we present a computationally efficient framework for jointly optimizing per-UE bandwidth allocation and UE–AP association assignment in the UL of a NS-enabled CF mMIMO system, aiming to maximize the weighted sum-rate while satisfying heterogeneous QoS requirements for eMBB and URLLC UEs. Owing to the NP-hard nature of the formulated problem, it is decomposed into two tractable sub-problems, each addressed by an efficient heuristic scheme designed to ensure fair QoS distribution even under minimum-bandwidth constraint violation scenarios. Numerical results demonstrate that the proposed scheme achieves a favorable balance between computational efficiency, weighted sum-rate performance, and QoS satisfaction, making it a robust choice for QoS-critical network deployments. While this work focuses on maintaining service for all UEs through fair QoS distribution under resource scarcity, future extensions could explore admission control strategies to further enhance strict QoS guarantees for latency-critical applications.
@@ A-A Proof of Theorem 1
Problem P0 contains $KM$ binary association variables $a_{k,m}\in\{0,1\}$ and $K$ continuous bandwidth allocation variables $b_{k}\in\mathbb{R}_{+}$, so it is a mixed-integer program. Nonconvexity follows from the objective’s dependence on the quadratic/bilinear terms (e.g., $(\sum_{m}a_{k,m}\gamma_{k,m})^{2}$ and $R_{k}=\left(1-\frac{\tau_{p}}{\tau_{c}}\right)b_{k}\log_{2}(1+\mathrm{SINR}_{k}(\mathbf{A}))$), together with the discrete feasible set $\{0,1\}^{KM}$.
For hardness, consider a restricted instance where bandwidths $b_{k}$ are fixed and QoS constraints are removed. Further, choose channel parameters (or assume negligible inter-UE interference) so that each $\mathrm{SINR}_{k}$ depends only on the AP assigned to UE $k$; then the objective reduces to maximizing $\sum_{k}w_{k}\log_{2}(1+\mathrm{SINR}_{k}(\mathbf{A}))$ which depends only on the assignment variables. Under this construction P0 subsumes the generalized assignment problem (GAP): assigning items (UEs) to bins (APs) with capacity constraints and profits, which is NP-hard [10]. Hence, by reduction from GAP, P0 is NP-hard.